OpenStudy (anonymous):

Personal music players are on sale for $39.75. This is $6.50 less than the regular price. Which equation could be used to find the regular price, r?

OpenStudy (anonymous):

The regular price, minus 6.50, will equal the sale price. So, if r is the regular price: r-6.50=39.75

OpenStudy (anonymous):

39.75-r=6.50 39.75+r=6.50 r-6.50=39.75 r+6.50=39.75